# SIM-Shapley: A Stable and Computationally Efficient Approach to Shapley Value Approximation

### v1

@@ 5.1 Benchmark Results on Standard ML Tasks
Figure 1: Comparison of convergence time (in seconds) on the Credit and Airbnb datasets. All methods incur a nonzero initialization time due to the computational cost of the first iteration. KernelSHAP may lack a valid convergence ratio during early iterations.
(a) Airbnb - Global - Consistency
(b) Airbnb - Global - Running Time
(c) Bank - Local - Consistency
(d) Bank - Local - Running Time
Figure 2: Boxplots of consistency (Pearson correlation) and running time (50 trials) for both global (Airbnb) and local (Bank) explainability settings. SIM-Shapley and Stable-SIM-Shapley are compared against baseline methods on both datasets.
5.2 Real-World Interpretability Assessment
To validate our methods in a realistic clinical setting, we apply SIM-Shapley and Stable-SIM-Shapley to the MIMIC-IV-ED dataset (johnson2023mimic, ), focusing on in-hospital mortality prediction (see Appendix D.3.1 for dataset details). In addition to quantitative results (Section 5.1), we incorporate expert review from a board-certified emergency physician, who confirms that the top-ranked features–such as vital signs, age, and comorbidity burden–are aligned with established medical knowledge. This supports the interpretability and practical utility of our methods. Full commentary appears in Appendix D.3.2.
Table 3: Top-10 features identified by each SV method on the MIMIC cohort. A check mark (✓) indicates inclusion in the method’s top-10 ranking. Results are shown for both global and local explanations.
@@ 5.2 Real-World Interpretability Assessment
SAGE SIM-Shapley Stable-SIM KernelSHAP SIM-Shapley Stable-SIM
@@
6 Discussion
We have shown that reframing SV estimation as a stochastic optimization task—combining mini-batch sampling with momentum-driven updates and $\ell_{2}$ regularization—yields substantial gains in both speed and stability. Our SIM-Shapley algorithm converges linearly to the true solution (Theorem 2), with provable variance reduction at each step (Theorem 1), and registers negligible bias (Section 3.2). Empirically, on both standard benchmarks and the MIMIC-IV-ED clinical cohort, SIM-Shapley and its Stable variant reduce compute time by up to 85% while producing explanations that align with domain knowledge.
Crucially, our approach preserves the original weighted least-squares formulation, so all existing KernelSHAP enhancements—unbiased $A$ estimators, paired-sampling schemes, alternative cooperative-game definitions, and uncertainty quantification—can be dropped seamlessly into SIM-Shapley. The addition of an $\ell_{2}$ penalty strengthens convexity and invertibility, incurring no extra algorithmic complexity.
More broadly, our stochastic-iterative framework generalizes to any sampling-dependent quadratic minimization, offering a unifying foundation for accelerating Monte Carlo–based approximation tasks. Future work could explore adaptive momentum scaling, richer variance-reduction strategies, and extensions to other cooperative-game objectives beyond SV.
Like many existing SV approximation techniques, our method also adopts the marginal distribution as a proxy for the true conditional distribution, which may introduce approximation error in practice. Future work could explore more accurate conditional estimators—such as those based on deep generative models or copula methods—to improve attribution reliability. Additionally, it would be interesting to investigate hybrid frameworks that directly incorporate conditional sampling into the SV estimation process while preserving computational efficiency.
Appendix A Complete Proofs
Remark 1.
Proof.
Recall that the matrix $A$ is defined as:
$A=\frac{1}{m}\sum_{i=1}^{m}z_{i}z_{i}^{\top},$
[…]
for diagonal entries ($k=\ell$):
$\mathbb{E}[z_{ik}^{2}]=\mathbb{E}[z_{ik}]=\frac{1}{2};$
for off-diagonal entries ($k\neq\ell$):
Thus, the population matrix takes the explicit form:
As a result, its eigenvalues are:
and $\alpha:=\text{min}(\lambda_{1},\lambda_{2})=\frac{1}{4}$. ∎
A.1 Poof of Theorem 1
Proof.
Recall the estimator of KernelSHAP in [16] is defined as $\bm{\beta}$ and the recursive form of $\bm{\beta}^{(n+1)}$ is as following:
$\bm{\beta}^{(n+1)}$ $=(1-t)\sum_{j=1}^{n+1}t^{n+1-j}\bm{\delta}^{(j)}.$
The variance propagates recursively as:
$\vdots$
thus the recursion remains stable if both variances are controlled.
A.2 Poof of Theorem 2
Proof.
A.2.1 Existence of a Fixed Point
Given a vector $\bm{\beta}$, define $\delta^{*}(\bm{\beta})$ as the unique solution to the population-level version of problem Eq. (7), where all empirical averages are replaced by their exact expectations. Define the mapping:
A vector $\bm{\beta}^{*}$ is a fixed point if
This implies directly that
$\delta^{*}(\bm{\beta}^{*})=\bm{\beta}^{*},$
Since the objective function in Eq. (7) is strictly convex and smooth, and the constraint is affine, the solution mapping $\delta^{*}(\bm{\beta})$ is continuous in $\bm{\beta}$, implying that $T$ is continuous. Moreover, due to the quadratic growth of the objective, there exists a compact, convex set $\mathcal{K}\subset\mathbb{R}^{d}$ (e.g., a sufficiently large closed ball) such that $T(\mathcal{K})\subset\mathcal{K}$.
A.2.2 Local Linearization and Convergence Rate via the Jacobian
Let the error at iteration $n$ be:
$e^{(n)}=\bm{\beta}^{(n)}-\bm{\beta}^{*}.$
where
$\delta^{*}(\bm{\beta})\approx\bm{\beta}^{*}+M(\bm{\beta}-\bm{\beta}^{*}).$
Substituting into the update rule:
$\bm{\beta}^{(n+1)}=t\,\bm{\beta}^{(n)}+(1-t)\,\delta^{*}(\bm{\beta}^{(n)}),$
we obtain:
Subtracting $\bm{\beta}^{*}$ from both sides yields:
$e^{(n+1)}\approx\left[tI+(1-t)M\right]e^{(n)}.$
Define the linear update operator $G:=tI+(1-t)M$. Then, the convergence behavior is governed by the spectral radius:
$\rho(G)=\max\{|\lambda|:\lambda\in\sigma(G)\}.$
If $\rho(G)<1$, the error converges linearly:
$\|e^{(n)}\|=O(\rho(G)^{n}).$
In our setting, the matrix $A$ is given by:
$A=\frac{1}{m}\sum_{i=1}^{m}z_{i}z_{i}^{T},$
Thus, the linear operator $G$ becomes:
$Gx\approx t(1-\mu)x.$
Hence, the spectral radius of $G$ is:
where $\alpha_{\min}^{+}$ denotes the smallest nonzero eigenvalue of $A$, which has a lower bound (Remark 1). Since $10<t<$ 1 and $\frac{\lambda}{\alpha_{i}+\lambda}<1$, we conclude $\rho(G)<1$, and the iteration converges linearly:
$\|e^{(n)}\|=O(\rho(G)^{n}).$
Appendix B SIM-Shapley: Global Explanation Version
We exhibit SIM-Shapley for global explanation in Algorithm 2.
Algorithm 2 SIM-Shapley (Global Explanation)
0:  Final estimate $\bm{\beta}^{(n)}$
2:  for $n=1$ to $T$ do
3:     Independently sample $\{z_{j}\}_{j=1}^{m}\sim p(z)$
4:     Sample $\{(x_{j},y_{j})\}_{j=1}^{B}\sim p(U)$
6:     Solve Eq. (7) to obtain $\bm{\delta}^{(n)}$ and update $\bm{\beta}^{(n)}$
10:     if convergence criterion satisfied then
11:        break
12:     end if
13:  end for
14:  return  $\bm{\beta}^{(n)}$
Appendix C Stable SIM-Shapley
New iteration form with correcting initial deviation:
[…]
New explicit solution with correcting initial deviation:
$c_{1}$ $=\frac{t}{1-t^{n+1}},\quad c_{2}=\frac{1-t}{1-t^{n+1}}$ (16)
Algorithm 3 Stable SIM-Shapley (Local Explanation)
0:  Final estimate $\bm{\beta}^{(n)}$
2:  for $n=1$ to $T$ do
3:     Independently sample $\{z_{j}\}_{j=1}^{m}\sim p(z)$
5:     Solve Eq. (15) to obtain $\bm{\delta}^{(n)}$ and update $\bm{\beta}^{(n)}$
9:     if convergence criterion satisfied then
10:        break
11:     end if
12:     if negative sampling detected then
14:     end if
15:  end for
16:  return  $\bm{\beta}^{(n)}$
Algorithm 4 Stable SIM-Shapley (Global Explanation)
0:  Final estimate $\bm{\beta}^{(n)}$
2:  for $n=1$ to $T$ do
3:     Independently sample $\{z_{j}\}_{j=1}^{m}\sim p(z)$
4:     Sample $\{(x_{j},y_{j})\}_{j=1}^{B}\sim p(U)$
6:     Solve Eq. (15) to obtain $\bm{\delta}^{(n)}$ and update $\bm{\beta}^{(n)}$
10:     if convergence criterion satisfied then
11:        break
12:     end if
13:     if negative sampling detected then
15:     end if
16:  end for
17:  return  $\bm{\beta}^{(n)}$
Appendix D Additional Experimental Details
D.1 Pearson’s Correlation Coefficient as a Consistency Metric
To evaluate this consistency, prior work has employed regression analysis. As illustrated in Figure 1, the estimates from different methods tend to align closely along a straight line, indicating a strong linear relationship. This observation motivates the use of Pearson’s correlation coefficient as a quantitative measure of consistency. A coefficient near  11\,1\,1 suggests near-perfect linear agreement between two sets of Shapley estimates.
D.2 ML Tasks Details
eTable 1: Dataset profile and experiment settings. $t$: the hyperparameter in update formula, which may be different in global and local explanation.
Dataset Total Size Feature Dimension Background Set Size Global Reference Set Size $\mathbf{z}$ Sample Size Batch Size $\mathbf{t}$ (Global/Local)
Bike 10886 12 512 1088 120 512 0.55 / 0.50
Credit 1000 20 256 100 200 512 0.55 / 0.55
Bank 45210 16 512 2000 160 512 0.55 / 0.50
Airbnb 48590 15 1024 4859 150 512 0.55 / 0.50
MIMIC-IV 414537 56 2048 10000 560 512 0.55 / 0.55
For baselines, we compare against SAGE algorithm proposed in [2] for global explanations, and original KernelSHAP [5] for local explanations, excluding the enhancements proposed in [16], which remain applicable to SIM-Shapley.
Figures 23 and 4 present additional results consistent with the main paper. Their descriptions follow the same conventions and support similar conclusions.
eTable 2: Comparison of two neural network architectures.
Layer # MLP for Airbnb MLP for MIMIC-IV
1 Linear(15, 512) Linear(56, 512)
2 ELU ELU
3 Linear(512, 512) Linear(512, 256)
4 ELU ELU
5 Linear(512, 1) Linear(256, 1)
6 Sigmoid
(a) Airbnb - Local - Consistency
(b) Airbnb - Local - Running Time
(c) Bike - Local - Consistency
(d) Bike - Local - Running Time
(e) Credit - Local - Consistency
(f) Credit - Local - Running Time
(a) Bank - Global - Consistency
(b) Bank - Global - Running Time
(c) Bike - Global - Consistency
(d) Bike - Global - Running Time
(e) Credit - Global - Consistency
(f) Credit - Global - Running Time
(a) Bank - Global Explanation
(b) Bank - Local Explanation
(c) Bike - Global Explanation
(d) Bike - Local Explanation
eFigure 4: Comparison of convergence time(s). Compare global and local interpretability methods on Bank and Bike datasets.
D.3 MIMIC-IV Experimental Details
D.3.1 Cohort Formation
The MIMIC-IV Emergency Department (MIMIC-IV-ED) dataset [37] is a publicly available resource containing over 400,000 emergency department (ED) visit episodes. Following the data extraction pipeline proposed by Xie et al. [38], we constructed a master dataset and formed our study cohort by excluding encounters with missing values in any of the 56 commonly used ED variables–such as demographics, vital signs, and comorbidities–selected for analysis.
D.3.2 Domain-Guided Interpretation of Results
To assess the clinical relevance of the feature attributions generated by SIM-Shapley and Stable-SIM-Shapley on the MIMIC-IV-ED dataset, we consulted a board-certified emergency physician. The physician independently reviewed the top-ranked features identified by our methods for predicting inpatient mortality and confirmed that they are plausible and aligned with clinical understanding. Notably, features such as vital signs, age, and comorbidity burden were consistent with known risk factors in emergency medicine. This alignment with expert knowledge supports the interpretability and practical utility of our proposed method in real-world clinical settings. These observations align with domain standards such as the Canadian Triage and Acuity Scale (CTAS) [^2] https://files.ontario.ca/moh_3/moh-manuals-prehospital-ctas-paramedic-guide-v2-0-en-2016-12-31.pdf, and comorbidity scoring systems like the Charlson Comorbidity Index [39, 40].

### v2

@@ 6 Sigmoid
(e) Bank - Global Explanation
(f) Bank - Local Explanation
(g) Bike - Global Explanation
(h) Bike - Local Explanation
eFigure 4: Comparison of convergence time(s). Compare global and local interpretability methods on standard machine learning datasets.
eTable 3: Average bias ($\ell_{2}$ distance) and runtime (seconds $\times 10^{3}$) across sample sizes on Credit dataset. Times for SIM-Shapley and KernelSHAP correspond to estimation time, whereas FastSHAP entries report the neural network training time. FastSHAP* denotes the reduced-parameter version.
Sample Size
Method Metric 64 256 512 768 1024 1536 2048 2304 2816 3840
SIM-Shapley Bias 0.028 0.028 0.019 0.019 0.019 0.018 0.018 0.018 0.018 0.018
Time 6.92 5.81 5.97 10.64 12.27 14.70 19.88 26.44 33.82 42.26
KernelSHAP Bias 0.012 0.012 0.012 0.009 0.008 0.007 0.006 0.006 0.005 0.004
Time 8.36 11.66 9.87 15.05 15.37 23.83 30.65 32.99 34.56 54.91
FastSHAP Bias 0.12
Time 12800
FastSHAP* Bias 0.13
Time 12190
E.4 Real-World Interpretability Assessment
To validate our methods in a realistic clinical setting, we apply SIM-Shapley and Stable-SIM-Shapley to the MIMIC-IV-ED dataset (Johnson et al., 2023), focusing on in-hospital mortality prediction (see Appendix E.4.1 for dataset details). In addition to quantitative results (Section 4.3), we incorporate expert review from a board-certified emergency physician, who confirms that the top-ranked features (provided in Table 4)–such as vital signs, age, and comorbidity burden–are aligned with established medical knowledge. This supports the interpretability and practical utility of our methods. Full commentary appears in Appendix E.4.2.
E.4.1 Cohort Formation
The MIMIC-IV Emergency Department (MIMIC-IV-ED) dataset (Johnson et al., 2023) is a publicly available resource containing over 400,000 emergency department (ED) visit episodes. Following the data extraction pipeline proposed by Xie et al. (2022), we constructed a master dataset and formed our study cohort by excluding encounters with missing values in any of the 56 commonly used ED variables–such as demographics, vital signs, and comorbidities–selected for analysis.
E.4.2 Domain-Guided Interpretation of Results
To assess the clinical relevance of the feature attributions generated by SIM-Shapley and Stable-SIM-Shapley on the MIMIC-IV-ED dataset, we consulted a board-certified emergency physician. The physician independently reviewed the top-ranked features identified by our methods for predicting inpatient mortality and confirmed that they are plausible and aligned with clinical understanding. Notably, features such as vital signs, age, and comorbidity burden were consistent with known risk factors in emergency medicine. This alignment with expert knowledge supports the interpretability and practical utility of our proposed method in real-world clinical settings. These observations align with domain standards such as the Canadian Triage and Acuity Scale (CTAS) [^3] https://files.ontario.ca/moh_3/moh-manuals-prehospital-ctas-paramedic-guide-v2-0-en-2016-12-31.pdf, and comorbidity scoring systems like the Charlson Comorbidity Index (Schuttevaer et al., 2022; Murray et al., 2006).
eTable 4: Top-10 features identified by each SV method on the MIMIC cohort. A check mark (✓) indicates inclusion in the method’s top‑10 ranking. Results are shown for both global and local explanations.
@@ E.4.2 Domain-Guided Interpretation of Results
SAGE SIM‑Shapley Stable‑SIM KernelSHAP SIM‑Shapley Stable‑SIM